To limit light pollution on to the dome floor a baffle was made around the top calibration-unit Halogen lamp vent. A baffle should also be made around the fiber-spectrograph coupling, which currently lets in stray light.

The electronics of the FIES camera focus mechanism that was 'affected' by the lightning in March 2007 still needs to be repaired and we should purchase some more spare parts.

One recurrent problem is the mounting of the fiber arm on the top calibration unit. This often becomes mis-aligned and needs some proper fixing.

One thing that still needs to be decided is whether to upgrade the old fibre bundle to serve as a proper spare.

An exposure meter was received from Aarhus and after some investigation a proper local was found and the meter was recently installed. The main task now is to test it, particular under different weather circumstances, after which it should be determined how it can be incorporated in our system. One of the objectives is to be able to use the results from the exposure meter directly to determine when a certain signal-to-noise is reached in an exposure, instead of having to guess the required exposure time which is very difficult if the weather is not optimal.

Thomas Augusteijn 2008-11-14